OpenStudy (muhammadturawa):

f(x) = sin2x g(x)= cos(30'-x) Determine algebraically the value of x for which f(x) = g(x)

OpenStudy (muhammadturawa):

pls help

OpenStudy (anonymous):

sin2x=cos(30-x) cos(30-x)=cos30cosx+sin30sinx

OpenStudy (anonymous):

so how did you solve

OpenStudy (anonymous):

?

OpenStudy (anonymous):

stop!

OpenStudy (anonymous):

i was wrong cosx=sin(90-x cos(30-x)=sin(90-(30-x))

OpenStudy (anonymous):

sin(60+x)

OpenStudy (anonymous):

sin2x=sin(60+x)

OpenStudy (anonymous):

do you see it?

OpenStudy (anonymous):

do you see how i got there?

OpenStudy (anonymous):

2X=60+X